Michael R. DeCorte

2300 Naudain St, Phily, PA, (215) 546-0497 mrd@sun.soe.clarkson.edu

Objective: Gee, why can’t I put “To play with cool toys and have fun”?

Education: Major: Bachelor of Science, Computer Science, Clarkson University May 1989

Publications: My picture was in my HS newpaper onc. Does that count?

Projects: Eclipse During a lunar eclipse the moon projects its shadow over the earth.

Experience:

Unix Wizard  September 1986 to Present

Math and Computer Science,Clarkson University— Well, people keep asking me all these nifty questions about Unix so that makes me a Wizard right?

TEX Wizard  September 1986 to Present

Math and Computer Science,Clarkson University— See Unix Wizard for a description and substitue TEX for Unix.

Systems:

• Unix 4.x  • Unix V  • Sun OS  • VMS

References: Call me I will see if I can dig someone up. I hope my old room mate counts.